# Error Broadcast and Decorrelation as a Potential Artificial and Natural Learning Mechanism

### v2

@@ I.3.7 Hyperparameters for 10-layer CorInfoMax-EBD on MNIST and CIFAR-10 datasets for batch size 1
Table 11: 10-Layer CorInfoMax-EBD hyperparameters: MNIST dataset, $1$-batch size.
@@
$B$ $1$
I.4 Multi-Layer Perceptron
In this section, we provide additional details about the numerical experiments conducted to train Multi-layer Perceptrons (MLPs) using the EBD algorithm (MLP-EBD). Appendix I.4.1 outlines the implementation details of these experiments, while Appendix I.4.2 discusses the initialization of algorithm variables. Information about hyperparameters and their values for the MNIST and CIFAR-10 datasets can be found in Appendices I.4.3-I.4.4. In Appendix I.2 we present the accuracy and loss learning curves for the MLP architecture, shown in Figures 5h.(a)-(b) and Figures 6.(a)-(b), respectively.
I.4.1 Implementation details
For the MLP experiments using the proposed EBD approach, we adopted the same network architecture as described in [9], detailed in Tables 3 and 4.
@@ I.4.1 Implementation details
I.4.2 Initialization of algorithm variables
@@ I.4.3 Description of hyperparameters
$WS$ Weight Sparsity
I.4.4 Hyperparameters for MNIST and CIFAR-10 datasets
Table 13 and 14 list the hyperparameters used in the MLP-EBD numerical experiments for the MNIST and CIFAR-10 datasets respectively, where $m$ denotes the iteration index in both.
Table 13: Hyperparameters for MLP-EBD for the MNIST dataset.
Hyperparameter Value
$\alpha[m]$ $\frac{1}{1.5\times\lfloor\frac{m}{10}\rfloor+1}$
$\alpha_{2}[m]$ ${\lfloor\frac{m}{10}\rfloor/3e4+1}$
$\mu^{(d,b,k)}[m]$ $18000\alpha[m]\alpha_{2}[m]$ for $k=0,1$
@@ I.4.4 Hyperparameters for MNIST and CIFAR-10 datasets
$\mu^{(d,f,k)}[m]$ $0.005\alpha[m]\alpha_{2}[m]$ for $k=0,1$
$\mu^{(E,k)}[m]$ $\left[\begin{array}[]{ccc}2.5e-4&1.5e-3&0\end{array}\right]\alpha[m]$
$\mu^{(p,k)}[m]$ $\left[\begin{array}[]{ccc}4e-3&6e-3&1e-10\end{array}\right]\alpha[m]$
$p^{(k)}[m]$ $\left[\begin{array}[]{ccc}2.5e-1&2.5e-1&0.1\end{array}\right]\alpha[m]$
$\mu^{(k)}_{\ell_{1}}$ $\left[\begin{array}[]{ccc}8e-1&3e-1&0\end{array}\right]\alpha[m]$
$\mu^{(k)}_{w-\ell_{2}}$ $1.6e-4\alpha[m]$ for all layers
$\lambda_{E}$ $0.99999$
$\lambda_{d}$ $0.999999$
$m^{(d)}$ $0.9999$ for all layers
$B$ 20
$WS$ $55$
Table 14: Hyperparameters for MLP-EBD for the CIFAR-10 dataset.
Hyperparameter Value
$\alpha[m]$ $\frac{1}{1.5\times\lfloor\frac{m}{10}\rfloor+1}$
$\alpha_{2}[m]$ ${\lfloor\frac{m}{10}\rfloor/3e4+1}$
$\mu^{(d,b,k)}[m]$ $\left[\begin{array}[]{cccc}4000&2000&2000&3500\end{array}\right]\alpha[m]\alpha_{2}[m]$
$\mu^{(d,f,k)}[m]$ $0.005\alpha[m]\alpha_{2}[m]$ for $k=0,1$
$\mu^{(E,k)}[m]$ $\left[\begin{array}[]{cccc}2.5e-4&1.5e-3&1.5e-3&0\end{array}\right]\alpha[m]$
$\mu^{(p,k)}[m]$ $\left[\begin{array}[]{cccc}4e-3&6e-3&6e-3&1e-10\end{array}\right]\alpha[m]$
$p^{(k)}[m]$ $\left[\begin{array}[]{cccc}2.5e-1&2.5e-1&2.5e-1&0.1\end{array}\right]\alpha[m]$
$\mu^{(k)}_{\ell_{1}}$ $\left[\begin{array}[]{cccc}8e-1&3e-1&3e-1&0\end{array}\right]\alpha[m]$
$\mu^{(k)}_{w-\ell_{2}}$ $1.6e-4\alpha[m]$ for all layers
$\lambda_{E}$ $0.99999$
$\lambda_{d}$ $0.999999$
$m^{(d)}$ $0.9999$ for all layers
$B$ 20
$WS$ $40$
I.5 Convolutional Neural Network
In this section, we offer additional details regarding the numerical experiments for training Convolutional Neural Neural Networks (CNNs) using EBD algorithm (CNN-EBD). Section I.5.1 provides information about implemetation details. Appendices I.5.2 and I.5.3 discuss the initialization of the algorithm’s variables and describe the hyperparameters. Finally, Appendix I.5.4 detail the specific hyperparameter configurations used in our numerical experiments for the MNIST and CIFAR-10 datasets. In Appendix I.2 we present the accuracy and loss learning curves for the CNN, shown in Figures 5h.(c)-(d) and Figures 6.(c)-(d), respectively.
I.5.1 Implementation details
The architectures we utilized for the CNN networks can be found in tables 3 and 4 respectively for the MNIST and CIFAR10 datasets. In the training, we used the Adam optimizer with hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$ [53]. Also, the model biases are not utilized. In the CNN-EBD implementation the following loss and regularization functions as detailed in section D.1 are used:
@@ I.5.1 Implementation details
I.5.2 Initialization of algorithm variables
@@ I.5.3 Description of hyperparameters
I.5.4 Hyperparameters for MNIST and CIFAR-10 datasets
@@
Table 16: Hyperparameters for CNN-EBD for both the MNIST and CIFAR-10 datasets, where $i$ denotes the epoch index.
Hyperparameter MNIST CIFAR-10
$\alpha[i]$ $10^{-4}\cdot 0.97^{-i}$ $10^{-4}\cdot 0.97^{-i}$
$\mu^{(d,b,k)[i]}$ $0.1\alpha[i]$ for $k=0,1,2,3$ $0.1\alpha[i]$ for $k=0,1,2,3$
$10\alpha[i]$ for $k=4$ $10\alpha[i]$ for $k=4$
$\mu^{(E,k)}[i]$ $\left[\begin{array}[]{ccccc}1&1&1&10&0\end{array}\right]10^{-7}\alpha[i]$ $\left[\begin{array}[]{ccccc}1&1&1&1&1\end{array}\right]10^{-6}\alpha[i]$
$\mu^{(k)}_{\ell_{1}}[i]$ $\left[\begin{array}[]{ccccc}1&1&1&10&0\end{array}\right]10^{-11}\alpha[i]$ $\left[\begin{array}[]{ccccc}1&1&1&10^{2}&0\end{array}\right]10^{-10}\alpha[i]$
$\sigma_{{\mathbf{W}}}$ $\sqrt{\frac{1}{6}}$ $\sqrt{\frac{1}{6}}$
$\sigma_{\mathbf{{\mathbf{R}}}_{lin}}$ $21e-$ 2 $21e-$ 2
$\sigma_{\mathbf{{\mathbf{R}}}_{conv}}$ $21e-$ 2 $21e-$ 2
$\lambda_{d}$ $0.99999$ $0.99999$
$\lambda_{E}$ $0.99999$ $0.99999$
$\lambda_{R}$ $22e-$ 2 $22e-$ 2
$\beta$ $81e-$ 8 $51e-$ 5
$\epsilon_{L}$ $81e-$ 8 $81e-$ 8
$\epsilon$ $51e-$ 5 $51e-$ 5
$B$ $16$ $16$
I.6 Locally Connected Network
In this section, we offer additional details regarding the numerical experiments for the training of Locally Connected Networks (LCs) using EBD algorithm (LC-EBD). Appendix I.6.1 provides information about implemetation details. Appendices I.6.2 and I.6.3 discuss the initialization of the algorithm’s variables and describe the hyperparameters. Finally, Appendix I.6.4 detail the specific hyperparameter configurations used in our numerical experiments for the MNIST and CIFAR-10 datasets. In Appendix I.2 we present the accuracy and loss learning curves for the LCs, shown in Figures 5h.(e)-(f) and Figures 6.(e)-(f), respectively.
I.6.1 Implementation details
@@
I.6.2 Initialization of algorithm variables
@@ I.6.3 Description of hyperparameters
I.6.4 Hyperparameters for MNIST and CIFAR-10 datasets
@@
$\alpha[i]$ $10^{-4}\cdot 0.96^{-i}$ $10^{-4}\cdot 0.98^{-i}$
@@
$\sigma_{\mathbf{{\mathbf{R}}}_{lin}}$ $1$ $31e-$ 3
$\sigma_{\mathbf{{\mathbf{R}}}_{local}}$ $1$ $11e-$ 1
@@
$\lambda_{R}$ $23e-$ 2 $23e-$ 2
$\beta$ $81e-$ 8 $61e-$ 6
$\epsilon_{L}$ $81e-$ 8 $81e-$ 8
$\epsilon$ $51e-$ 5 $51e-$ 5
@@
I.7 Implementation details for Direct Feedback Alignment (DFA) and backpropagation training
@@
For BP-based models trained on MNIST and CIFAR-10, we used the Adam optimizer with hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$ [53]. For DFA and DFA+E models, we again used the Adam optimizer for CNN and LC models, while MLP models were trained using SGD with momentum.
In Tables 18 and 19, we detail the hyperparameters for models trained with BP, DFA, and DFA+E update rules on MNIST and CIFAR-10 respectively. Some of the learning rate and the learning rate decay values or methodologies are linked to the tables corresponding to the hyperparameter details of its EBD counterpart, where the same method is also utilized for its DFA or DFA+E counterpart. Unlinked values denote a constant value applied to each layer, or the step decay multiplier applied per epoch. Additionally, sparsity inducing losses are not utilized for BP, DFA and DFA+E models.
@@
Model Method Learning Rate ($\mu^{(d,b,k)}$) L2 Reg. Coef. LR Decay Epochs
BP $55e-$ 5 $51e-$ 5 0.96 120
@@
BP $55e-$ 5 $81e-$ 8 0.97 100
CNN DFA Table-16 $81e-$ 8 0.97 100
DFA+E Table-16 $81e-$ 8 0.97 100
BP $55e-$ 5 $81e-$ 8 0.96 100
LC DFA Table-17 $81e-$ 8 0.96 100
DFA+E Table-17 $81e-$ 8 0.96 100
@@
Model Method Learning Rate ($\mu^{(d,b,k)}$) L2 Reg. Coef. LR Decay Epochs
BP $55e-$ 5 $51e-$ 5 0.85 120
MLP DFA Table-14 0 Table-14 120
DFA+E Table-14 0 Table-14 120
BP $55e-$ 5 $51e-$ 5 0.92 200
CNN DFA Table-16 $51e-$ 5 0.97 200
DFA+E Table-16 $51e-$ 5 0.97 200
BP $41e-$ 4 $61e-$ 6 0.90 200
LC DFA Table-17 $61e-$ 6 0.96 200
DFA+E Table-17 $61e-$ 6 0.96 200
I.8 Runtime comparisons for the update rules

### v3

@@ I.2.6 Hyperparameters for 10-layer CorInfoMax-EBD on MNIST and CIFAR-10 datasets for batch size 1
Table 11: 10-Layer CorInfoMax-EBD hyperparameters: MNIST and CIFAR-10 datasets ($B=1$).
@@
$B$ 11
I.3 Multi-Layer Perceptron
In this section, we provide additional details about the numerical experiments conducted to train Multi-layer Perceptrons (MLPs) using the EBD algorithm (MLP-EBD). Appendix I.3.1 outlines the implementation details of these experiments, while Appendix I.3.2 discusses the initialization of algorithm variables. Information about hyperparameters and their values for the MNIST and CIFAR-10 datasets can be found in Appendices I.3.3-I.3.4. In Appendix I.10 we present the accuracy and loss learning curves for the MLP architecture, shown in Figures 6.(a)-(b) and Figures 7.(a)-(b), respectively.
I.3.1 Implementation details
For the MLP experiments using the proposed EBD approach, we adopted the same network architecture as described in [9], detailed in Tables 4 and 5.
@@ I.3.1 Implementation details
I.3.2 Initialization of algorithm variables
@@ I.3.3 Description of hyperparameters
WSWS Weight Sparsity
I.3.4 Hyperparameters for MLP-EBD on MNIST and CIFAR-10 Datasets
Table 13 summarizes the hyperparameters used in the MLP-EBD experiments for the MNIST and CIFAR-10 datasets. The iteration index is denoted by $m$ in all expressions.
Table 13: MLP-EBD hyperparameters for MNIST and CIFAR-10 datasets.
Hyperparameter MNIST CIFAR-10
$\alpha[m]$ $\dfrac{1}{1.5\times\lfloor\tfrac{m}{10}\rfloor+1}$ $\dfrac{1}{1.5\times\lfloor\tfrac{m}{10}\rfloor+1}$
$\alpha_{2}[m]$ $\dfrac{\lfloor\tfrac{m}{10}\rfloor}{3\times 10^{4}}+1$ $\dfrac{\lfloor\tfrac{m}{10}\rfloor}{3\times 10^{4}}+1$
$\mu^{(d,b,k)}[m]$ $18000\,\alpha[m]\alpha_{2}[m]$ for $k=0,1$ $4000,\,2000,\,2000,\,3500]\,\alpha[m]\alpha_{2}[m]$
@@ I.3.4 Hyperparameters for MLP-EBD on MNIST and CIFAR-10 Datasets
$\mu^{(d,f,k)}[m]$ $0.005\,\alpha[m]\alpha_{2}[m]$ for $k=0,1$ $0.005\,\alpha[m]\alpha_{2}[m]$ for $k=0,1$
$\mu^{(E,k)}[m]$ $2.5\times 10^{-4},\,1.5\times 10^{-3},\,0]\,\alpha[m] [$$2.5\times 10^{-4},\,1.5\times 10^{-3},\,1.5\times 10^{-3},\,0]\,\alpha[m]$
$\mu^{(p,k)}[m]$ $4\times 10^{-3},\,6\times 10^{-3},\,1\times 10^{-10}]\,\alpha[m] [$$4\times 10^{-3},\,6\times 10^{-3},\,6\times 10^{-3},\,1\times 10^{-10}]\,\alpha[m]$
$p^{(k)}[m]$ [0.25, 0.25, $0.1]\,\alpha[m] [0.25, 0.25,$ $0.25,\,0.25,\,0.1]\,\alpha[m$]
[…]
$\mu^{(k)}_{w-\ell_{2}}$ $1.6\times 10^{-4}\,\alpha[m]$ for all layers $1.6\times 10^{-4}\,\alpha[m]$ for all layers
$\lambda_{E}$ $0.99999$ $0.99999$
$\lambda_{d}$ $0.999999$ $0.999999$
$m^{(d)}$ $0.9999$ for all layers $0.9999$ for all layers
$B$ 2020 2020
WSWS $55$ $40$
I.4 Convolutional Neural Network
In this section, we offer additional details regarding the numerical experiments for training Convolutional Neural Neural Networks (CNNs) using EBD algorithm (CNN-EBD). Section I.4.1 provides information about implemetation details. Appendices I.4.2 and I.4.3 discuss the initialization of the algorithm’s variables and describe the hyperparameters. Finally, Appendix I.4.4 detail the specific hyperparameter configurations used in our numerical experiments for the MNIST, CIFAR-10 and CIFAR-100 datasets. In Appendix I.10 we present the accuracy and loss learning curves for the CNN, shown in Figures 6.(c)-(d) and Figures 7.(c)-(d), respectively.
I.4.1 Implementation details
The architectures we utilized for the CNN networks can be found in tables 4 and 5 respectively for the MNIST and CIFAR10 datasets. In the training, we used the Adam optimizer with hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$ [57]. Also, the model biases are not utilized. In the CNN-EBD implementation the following loss and regularization functions as detailed in section D.1 are used:
@@ I.4.1 Implementation details
Specifically for the CIFAR-100 experiments, we applied both training and test time augmentation of data to improve model generalization and handle increased difficulty in the task. At training time, each image was randomly translated into 2-pixels with reflection padding and a deterministic alternating horizontal flip that ensures that every image is flipped every other epoch, reducing redundancy compared to standard random flipping. During evaluation, we used test-time augmentation combining horizontal flipping and multi-crop averaging over six translated views (the original, two one-pixel translations, and their mirrored counterparts).
I.4.2 Initialization of algorithm variables
@@ I.4.3 Description of hyperparameters
$\alpha_{\text{exp}}$ Exponential learning rate decay parameter.
@@
I.4.4 Hyperparameters for MNIST, CIFAR-10 and CIFAR-100 datasets
@@
Table 15: Hyperparameters for CNN-EBD for the MNIST, CIFAR-10 and CIFAR-100 datasets, where $i$ denotes the epoch index.
Hyperparameter MNIST CIFAR-10 CIFAR-100
$\alpha_{\text{exp}}$ 0.97 0.97 0.97
$\alpha[i]$ $10^{-4}\cdot\alpha_{\text{exp}}^{-i}$ $10^{-4}\cdot\alpha_{\text{exp}}^{-i}$ $10^{-4}\cdot\alpha_{\text{exp}}^{-i}$
$\mu^{(d,b,k)[i]}$ $0.1\alpha[i]$ for $k=0,1,2,3$ $0.1\alpha[i]$ for $k=0,1,2,3$ $0.1\alpha[i]$ for $k=0,1,2,3,4$
$10\alpha[i]$ for $k=4$ $10\alpha[i]$ for $k=4$ $10\alpha[i]$ for $k=5$
$\mu^{(E,k)}[i]$ $\left[\begin{array}[]{ccccc}1&1&1&10&0\end{array}\right]10^{-7}\alpha[i]$ $\left[\begin{array}[]{ccccc}1&1&1&1&1\end{array}\right]10^{-6}\alpha[i]$ $\left[\begin{array}[]{cccccc}0&0&0&5&5&5\end{array}\right]10^{-7}\alpha[i]$
$\mu^{(k)}_{\ell_{1}}[i]$ $\left[\begin{array}[]{ccccc}1&1&1&10&0\end{array}\right]10^{-11}\alpha[i]$ $\left[\begin{array}[]{ccccc}1&1&1&10^{2}&0\end{array}\right]10^{-10}\alpha[i]$ $\left[\begin{array}[]{cccccc}0&0&0&1&1&0\end{array}\right]10^{-7}\alpha[i]$
$\sigma_{{\mathbf{W}}}$ $\sqrt{\frac{1}{6}}$ $\sqrt{\frac{1}{6}}$ $\sqrt{\frac{1}{6}}$
$\sigma_{\mathbf{{\mathbf{R}}}_{lin}}$ 1e21e-2 1e21e-2 1e21e-2
$\sigma_{\mathbf{{\mathbf{R}}}_{conv}}$ 1e21e-2 1e21e-2 1e21e-2
$\lambda_{d}$ $0.99999$ $0.99999$ $0.99999$
$\lambda_{E}$ $0.99999$ $0.99999$ $0.99999$
$\lambda_{R}$ 2e22e-2 2e22e-2 2e22e-2
$\beta$ 1e81e-8 1e51e-5 1e51e-5
$\epsilon_{L}$ 1e81e-8 1e81e-8 1e81e-8
$\epsilon$ 1e51e-5 1e51e-5 1e51e-5
$B$ $16$ $16$ $16$
$p$ N/A N/A $0.075$
I.5 Locally Connected Network
In this section, we offer additional details regarding the numerical experiments for the training of Locally Connected Networks (LCs) using EBD algorithm (LC-EBD). Appendix I.5.1 provides information about implemetation details. Appendices I.5.2 and I.5.3 discuss the initialization of the algorithm’s variables and describe the hyperparameters. Finally, Appendix I.5.4 detail the specific hyperparameter configurations used in our numerical experiments for the MNIST and CIFAR-10 datasets. In Appendix I.10 we present the accuracy and loss learning curves for the LCs, shown in Figures 6.(e)-(f) and Figures 7.(e)-(f), respectively.
I.5.1 Implementation details
@@
I.5.2 Initialization of algorithm variables
@@ I.5.3 Description of hyperparameters
I.5.4 Hyperparameters for MNIST and CIFAR-10 datasets
@@
$\alpha_{\text{exp}}$ 0.96 0.98
$\alpha[i]$ $10^{-4}\cdot\alpha_{\text{exp}}^{-i}$ $10^{-4}\cdot\alpha_{\text{exp}}^{-i}$
@@
$\sigma_{\mathbf{{\mathbf{R}}}_{lin}}$ 11 1e31e-3
$\sigma_{\mathbf{{\mathbf{R}}}_{local}}$ 11 1e11e-1
@@
$\lambda_{R}$ 3e23e-2 3e23e-2
$\beta$ 1e81e-8 1e61e-6
$\epsilon_{L}$ 1e81e-8 1e81e-8
$\epsilon$ 1e51e-5 1e51e-5
@@
$p$ N/A N/A
I.6 Implementation details for Direct Feedback Alignment (DFA) and backpropagation training
@@
For BP-based models trained on MNIST, CIFAR-10 and CIFAR-100, we used the Adam optimizer with hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$ [57]. For DFA and DFA+E models, we again used the Adam optimizer for CNN and LC models, while MLP models were trained using SGD with momentum.
In Tables 17 and 18, we detail the hyperparameters for models trained with BP, DFA, and DFA+E update rules on MNIST and CIFAR-10 respectively. In Table 19, we give the hyperparameters for the CNN model (detailed in Table 7) trained with with BP and DFA. Some of the learning rate and the learning rate decay values or methodologies are linked to the tables corresponding to the hyperparameter details of its EBD counterpart, where the same method is also utilized for its DFA or DFA+E counterpart. Unlinked values denote a constant value applied to each layer, or the step decay multiplier applied per epoch. Additionally, sparsity inducing losses are not utilized for BP, DFA and DFA+E models.
@@
Model Method Learning Rate ($\mu^{(d,b,k)}$) L2 Reg. Coef. LR Decay ($\alpha_{\text{exp}}$) Epochs
BP 5e55e-5 1e51e-5 0.96 120
@@
BP 5e55e-5 1e81e-8 0.97 100
CNN DFA Table-15 1e81e-8 0.97 100
DFA+E Table-15 1e81e-8 0.97 100
BP 5e55e-5 1e81e-8 0.96 100
LC DFA Table-16 1e81e-8 0.96 100
DFA+E Table-16 1e81e-8 0.96 100
@@
Model Method Learning Rate ($\mu^{(d,b,k)}$) L2 Reg. Coef. LR Decay ($\alpha_{\text{exp}}$) Epochs
BP 5e55e-5 1e51e-5 0.85 120
MLP DFA Table-13 0 Table-13 120
DFA+E Table-13 0 Table-13 120
BP 5e55e-5 1e51e-5 0.92 200
CNN DFA Table-15 1e51e-5 0.97 200
DFA+E Table-15 1e51e-5 0.97 200
BP 1e41e-4 1e61e-6 0.90 200
LC DFA Table-16 1e61e-6 0.96 200
DFA+E Table-16 1e61e-6 0.96 200
Table 19: Hyperparameter details for the CNN model trained on the CIFAR-100 dataset, including learning rate, L2 regularization coefficient, learning rate decay, dropout probability and number of epochs using BP and DFA methods.
Model Method Learning Rate ($\mu^{(d,b,k)}$) L2 Reg. Coef. LR Decay ($\alpha_{\text{exp}}$) Dropout ($p$) Epochs
CNN BP 5e55e-5 1e51e-5 0.97 0.5 200
DFA $\left[0.1\hskip 5.0pt0.1\hskip 5.0pt0.1\hskip 5.0pt1\hskip 5.0pt1\hskip 5.0pt50\right]\cdot 0.25\alpha[i]$ 0 0.95 0.075 300
I.7 Runtime comparisons for the update rules